I have to solve by the addition method. I am lost on this one because even when I rewrite the equation I still cant see it.
2x=2-y
3x+y=-1

Answer by checkley77(12844)

2x=2-y
2x+y=2
3x+y=-1 subtract
----------------------
-x=3
x=-3 ans.
2*-3=2-y
-6=2-y
y=2+6
y=8 ans.
Proof:
3*-3+8=-1
-9+8=-1
-1=-1